OpenStudy (anonymous):

for every percussion instrument in the band, there are 4 wind instruments in the band. if there are 68 wind instruments, how many percussion instruments are there? set up a proportion to solve!

OpenStudy (anonymous):

1/4=x/68 4x=68 x=17

OpenStudy (anonymous):

17 percussion instruments

OpenStudy (anonymous):

thank u